OpenStudy (anonymous):

Two cards are drawn at random from well shuffled deck of 52 cards. What is the probability that at least one of them is red?

hartnn (hartnn):

how many red cards are there?

OpenStudy (anonymous):

26

hartnn (hartnn):

so whats the probab. of drawing 1 red card from pack ?

OpenStudy (anonymous):

1/1326

OpenStudy (anonymous):



hartnn (hartnn):

probability of atleast 1 = 1- probability of none = 1- probability of NOT red *NOT red.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

By enumeration: Both red, first choice 26 of 52, second choice 25 of 51 = \[\frac{26}{52} * \frac{25}{51} = \frac{26}{2*51}\] First one or second one red, first choice 26 of 52, second choice 26 of 51 = \[2 * \frac{26}{52} * \frac{26}{51} = \frac{26}{51}\] Either one red = sum of the above = \[\frac{26}{2*51} + \frac{2*26}{2*51} = \frac{3*26}{2*51}\] I think.

OpenStudy (zarkon):

you have a typo \[\frac{26}{52} \cdot \frac{25}{51} =\frac{1}{2}\cdot\frac{25}{51}= \frac{25}{2\cdot51}\]